OpenStudy (blackops2_143):

what is k-2/12 divided by k when k = −4?

rebeccaxhawaii (rebeccaxhawaii):

substitute for k into the equation so -4-2/12 so -6/12

rebeccaxhawaii (rebeccaxhawaii):

then simplify -6/12=-1/2

OpenStudy (blackops2_143):

ok

OpenStudy (mathmale):

Please verify that we understand your expression correctly. Did you mean:\[\frac{ k-\frac{ 2 }{ 12 } }{ k }?\]

OpenStudy (blackops2_143):

no

OpenStudy (blackops2_143):



OpenStudy (blackops2_143):

and k=4

OpenStudy (mathmale):